# TiM4Rec: An Efficient Sequential Recommendation Model Based on Time-Aware Structured State Space Duality Model

### v2

@@ 4.1.4. Implementation Details
Our experimental procedures adhere to the paradigm requirements set by the PyTorch (paszke2019pytorch) and RecBole (recbole[1.2.0]). All models in the experiment are implemented with a model dimension of 64, which is the conventional choice for sequential recommendation models based on pure ID modeling. For models based on the Mamba architecture, such as Mamba4Rec, SSD4Rec* and TiM4Rec, the SSM state factor was consistently set to 32, the kernel size for 1D causal convolution is 4, and the block expansion factor for linear projections is 2. It is notable that the 1D causal convolutional layer within the Mamba architecture constrains the multi-head parameter to multiples of 8. However, in the context of sequential recommendation, this parameter typically holds values of 2 or 4. Consequently, we modify the 1D causal convolutional layer to accommodate the specific modeling tasks of sequential recommendation, ultimately standardizing it to a value of 4. The backbone layers across all models are consistently configured to a depth of 2. To address the sparsity of the Amazon datasets, a dropout rate of 0.4 is used, compared to 0.2 for MovieLens-1M and KuaiRand. The Adam optimizer (kingma2014adam) is used uniformly with a learning rate set to 0.01. A batch size of 2048 is employed for training, and the validated batch size is 4096. The maximum sequence length is set proportionally to the mean number of actions per user: 200 for MovieLens1M and 50 for Amazon-Beauty and KuaiRand datasets. We adhere to the optimal parameter configurations specific to each baseline, ensuring that the core parameters remain consistent across all models, and make adjustments to maximize performance to the extent feasible.
@@ 4.2. Performance Comparison
  • In low-dimensional settings, the SSD architecture experiences a considerable decline in recommendation performance compared to the SSM architecture. However, when compared with models founded on Transformer architectures, it exhibits enhanced performance. This observation aligns with our theoretical analysis presented in Section 3.3.
  • TiM4Rec exhibits a significant performance enhancement when compared to SSD4Rec*, which is modeled on a purely SSD-based architecture, thereby confirming the efficacy of the Time-aware SSD Block. Furthermore, TiM4Rec demonstrates a slight performance advantage over Mamba4Rec. This observation underscores the effectiveness of leveraging temporal differences in user interactions to capture shifts in user interests, effectively mitigating the performance decline observed when transitioning from the SSM architecture to the SSD architecture.
@@ 4.3. Dimension comparison
To assess the suitability of modeling pure ID sequence recommendation tasks in lower dimensions, we compare the experimental results presented in the SSD4Rec paper at 256 dimensions. These results are shown in Table 4. Our research findings clearly demonstrate that the Mamba4Rec model tends to exhibit superior performance in 64-dimensional space. Additionally, the SSD4Rec* model that we replicate, which does not utilize variable-length sequences or bidirectional SSDs, showed only a slight decline in performance at 64 dimensions compared to the results presented in the original SSD4Rec paper at 256 dimensions. Notably, the performance of SSD4Rec at 256 dimensions is even inferior to that Mamba4Rec at 64 dimensions. These findings provide robust evidence that pure ID sequence recommendation tasks are more effectively modeled in lower-dimensional spaces. Moreover, the computational overhead associated with 256 dimensions is substantially higher, underscoring the advantages of utilizing our model. In addtional, a key advantage of the SSD architecture over the SSM architecture is its superior capability in high-dimensional modeling. To validate this advantage, we conduct a comparative analysis of the performance of our proposed method, TiM4Rec, at a dimensionality of 256. The results confirm that TiM4Rec effectively preserves the benefits of the SSD architecture even in high-dimensional settings.
@@ 4.5. Ablation Study
To validate the efficacy of our proposed time-aware algorithm along with other critical design elements, we perform a series of ablation studies using the ML-1M dataset. The procedure for the systematic elimination of each component is detailed as follows:
@@
Table 6. Results of ablation studies on the ML-1M dataset
Component R@10 R@20 N@10 N@20 M@10 M@20
Full 0.3310 0.4338 0.1932 0.2194 0.1512 0.1584
w/o Time 0.3199 0.4240 0.1841 0.2104 0.1425 0.1498
w/o FFN 0.3247 0.4278 0.1842 0.2103 0.1415 0.1487
1 layer 0.3088 0.4146 0.1773 0.2040 0.1373 0.1446
3 layer 0.3310 0.4323 0.1916 0.2172 0.1490 0.1560

### v3

@@ 4.1.4. Implementation Details
Our experimental procedures adhere to the paradigm requirements set by the PyTorch (Pytorch2019) and RecBole (Recbole-1.2.0). All models in the experiment are implemented with a model dimension of 64, which is the conventional choice for sequential recommendation models based on pure ID modeling. For models based on the Mamba architecture, such as Mamba4Rec, SSD4Rec*, and TiM4Rec, the SSM state factor was consistently set to 32, the kernel size for 1D causal convolution was 4, and the block expansion factor for linear projections was 2. It is notable that the 1D causal convolutional layer within the Mamba architecture constrains the multi-head parameter to multiples of 8. However, in the context of sequential recommendation, this parameter typically holds values of 2 or 4 (SASRec2018; BERT4Rec2019). Consequently, we modify the 1D causal convolutional layer to accommodate the specific modeling tasks of sequential recommendation, ultimately standardizing it to a value of 4. The backbone layers across all models are consistently configured to a depth of 2. To address the sparsity of the Amazon datasets, a dropout rate of 0.4 is used, compared to 0.2 for MovieLens-1M and KuaiRand. The Adam optimizer (Adam2015) is used uniformly with a learning rate set to 0.01. A batch size of 2048 is employed for training, and the validated batch size is 4096. The maximum sequence length is set proportionally to the mean number of actions per user: 200 for MovieLens1M and 50 for Amazon-Beauty and KuaiRand datasets. We adhere to the optimal parameter configurations specific to each baseline, ensuring that the core parameters remain consistent across all models, and make adjustments to maximize performance to the extent feasible.
@@ 4.2. Performance Comparison
  • In low-dimensional settings, the SSD architecture experiences a considerable decline in recommendation performance compared to the SSM architecture. However, when compared with models founded on Transformer architectures, it exhibits enhanced performance. This observation aligns with our theoretical analysis presented in Section 3.3.1.
  • TiM4Rec exhibits a significant performance enhancement when compared to SSD4Rec*, which is modeled on a purely SSD-based architecture, thereby confirming the efficacy of the Time-aware SSD Block. Furthermore, TiM4Rec demonstrates a slight performance advantage over Mamba4Rec. This observation underscores the effectiveness of leveraging time-difference in user interactions to capture shifts in user interests, effectively mitigating the performance decline observed when transitioning from the SSM architecture to the SSD architecture.
@@ 4.3. Dimension comparison
To assess the suitability of modeling pure ID sequential recommendation tasks in lower dimensions, we compare the experimental results presented in the SSD4Rec paper at 256 dimensions. These results are shown in Table 4. Our research findings clearly demonstrate that the Mamba4Rec model tends to exhibit superior performance in 64-dimensional space. Additionally, the SSD4Rec* model that we replicate, which does not utilize variable-length sequences or bidirectional SSDs, showed only a slight decline in performance at 64 dimensions compared to the results presented in the original SSD4Rec paper at 256 dimensions. Notably, the performance of SSD4Rec at 256 dimensions is even inferior to that Mamba4Rec at 64 dimensions. These findings provide robust evidence that pure ID sequential recommendation tasks are more effectively modeled in lower-dimensional spaces. Moreover, the computational overhead associated with 256 dimensions is substantially higher, underscoring the advantages of utilizing our model. In addtional, a key advantage of the SSD architecture over the SSM architecture is its superior capability in high-dimensional modeling. To validate this advantage, we conduct a comparative analysis of the performance of our proposed method, TiM4Rec, at a dimensionality of 256. The results confirm that TiM4Rec effectively preserves the benefits of the SSD architecture even in high-dimensional settings.
@@ 4.4. Efficiency Analysis
Additionally, we compared the training and inference times of various models across different sequence lengths, as shown in Figure 6. The results indicate that TiM4Rec consistently maintains a significant speed advantage across all sequence lengths. Compared to the Transformer-based SASRec and the SSM-based Mamba4Rec, TiM4Rec effectively inherits the efficiency advantages of the SSD architecture while achieving superior performance. This result also demonstrates that our proposed time-aware enhancement method has linear computational complexity, ensuring that TiM4Rec does not experience efficiency degradation as sequence length increases.
(a) Training time comparison
(b) Inference time comparison
Figure 6. Comparison of model training and inference times across different sequence lengths on ML-1M. The results for SASRec and LRURec at a sequence length of 400 were obtained by reducing the batch size due to Out-Of-Memory.
@@ 4.5. Ablation Study
To validate the efficacy of our proposed time-aware algorithm along with other critical design elements, we perform a series of ablation studies using the ML-1M and Amazon-Beauty dataset. The procedure for the systematic elimination of each component is detailed as follows:
@@
Table 6. Results of ablation studies
Dataset $\rightarrow$ ML-1M Amazon-Beauty
Component R@10 R@20 N@10 N@20 M@10 M@20 R@10 R@20 N@10 N@20 M@10 M@20
Full 0.3310 0.4338 0.1932 0.2194 0.1512 0.1584 0.0854 0.1204 0.0446 0.0533 0.0321 0.0345
w/o Time 0.3199 0.4240 0.1841 0.2104 0.1425 0.1498 0.0806 0.1146 0.0423 0.0509 0.0306 0.0329
w/o FFN 0.3247 0.4278 0.1842 0.2103 0.1415 0.1487 0.0827 0.1183 0.0422 0.0512 0.0298 0.0322
1 layer 0.3088 0.4146 0.1773 0.2040 0.1373 0.1446 0.0739 0.1042 0.0437 0.0513 0.0344 0.0365
3 layer 0.3310 0.4323 0.1916 0.2172 0.1490 0.1560 0.0848 0.1211 0.0431 0.0522 0.0303 0.0328